What is the formula for solving this: Sour balls are 2 cents each, and lollipops are 5 cents each. Sally spent $1 and got 35 pieces of candy. How many sour balls did she get?
Answer by scott8148(6628)

two unknowns, two equations

"35 pieces of candy" ___ S + L = 35 ___ 2S + 2L = 70

"Sally spent $1" ___ 2S + 5L = 100

subtracting eqns ___ 3L = 30 ___ L = 10

substituting ___ S + 10 = 35 ___ S = 25